Let a=sin7+cos7,b=√sin7+√cos7,c=√1+sin14,d=1 then decreasing order of value is
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
a = sin 7 + cos7
b = sqrt(sin7) + sqrt (cos 7)
c = sqrt (1 + sin 14) = sqrt ( sin² 7 + cos² 7 + 2sin 7 cos7) = |sin 7 + cos 7| = sin 7 + cos7
d = 1
 
Hence 
 
b > a = c > d